OpenStudy (anonymous):

what is the slope of a line that passes through the origin and the point (6,3)

OpenStudy (anonymous):

m=1/2, The slope m of the line through the points (x1, y1) and (x 2, y 2) is given by. Slope Formula. you have two sets of points, (0,0) and (6,3) /the formula is (y2-y1)/(x2-x1) and (3-0)/(6-0)=3/6=1/2